Solve the system by substitution.
3x – y = –7
x + y = –9

Answer by malakumar_kos@yahoo.com(315)


Solve the system by substitution.
3x – y = –7
x + y = –9


x+y = -9 or y = -9-x........(1)
3x - y = -7........(2)
substituting for y from (1) in (2)
3x - (-9-x) = -7
3x+9+x = -7
4x = -7-9 = -16
x = -16/4 = -4
substituting for x in (1) we get y = -9-(-4) = -9+4 = -5
The required solution is x = -4, & y = -5